OpenStudy (anonymous):

Find x 4(x - 6) = -2(x - 3) I got x = 9 but the answers would be 12 = -12 and that doesn't work. Is this equation even possible for the answers to be equal?

OpenStudy (vijay):

4(x - 6) = -2(x - 3) 4x-24=-2x+6 6x=30 x=5